solve the following equation q: 2q+3/3+3q-4/6=q-2/2
Answer by mananth(16946)

2q+3/3 + 3q-4/6= q-2/2
.
2q+3/3 + 3q-4/6- q-2/2 =0
LCM =6
2(2q+3) + (3q-4) - 3(q-2) / 6=0
4q+6+3q-4-3q+6 =0
4q-8=0
4q=8
q=2